OpenStudy (anonymous):

find the critical points of the function: f(x) = 6x3 + x2 + 6x

OpenStudy (anonymous):

-18/8

OpenStudy (aum):

Find f'(x). Set it to zero and solve for x.

OpenStudy (anonymous):

it says that the answer you gave me is incorrect!

OpenStudy (anonymous):

i tried that but i can't seem to get the critical points because the factoring doesn't work out

OpenStudy (aum):

f(x) = 6x^3 + x^2 + 6x f'(x) = ?

OpenStudy (anonymous):

f'(x)=18x^2+2x+6

OpenStudy (aum):

18x^2 + 2x + 6 = 0 or 9x^2 + x + 3 = 0 solve for x.

OpenStudy (anonymous):

thats where i get stuck. it doesn't factor out evenly

OpenStudy (anonymous):

i don't know how to solve at that point

OpenStudy (aum):

Use the quadratic formula. If there are no real solutions there are no critical points.

OpenStudy (anonymous):

ok ill try that! thanks so much